Develop a quadratic function that fits the data in the table:


X --- | --- Y
3.99 | 500
5.99 | 650
7.99 | 200


I get how to do it in a larger sequence/table, but as this is only a set of 3 I don't really understand how to make a function, any help?

Answer by Alan3354(69443)

Develop a quadratic function that fits the data in the table:


X --- | --- Y
3.99 | 500
5.99 | 650
7.99 | 200
======================
If you know it's a quadratic 3 points is sufficient.
---
y = ax^2 + bx + c
Sub the values to get 3 equations
500 = a^3.99^2 + 3.99x + c
15.9201a + 3.99b + c = 500
---
35.8801a + 5.99b + c = 650
---
63.8401a + 7.99b + c = 200
========
Solve for a, b & c
Can you do that?
email via the TY note for help or to check your work.